# Hard rigid rods on Husimi trees

### v1

@@ III Results for rods on the square HL
This is indeed confirmed in Fig. 4, where the order parameter is shown as a function of the density of sites occupied by the rods on the lattice, for different values of $k$. As $k$ grows, $\Psi$ approaches unity in the limit of full occupancy (corresponding to $z\rightarrow\infty$ and, thus, $\rho=1$), but it does not reach it, as already demonstrated in Ref. rso22 . Therefore, in such a limit the stable phase is nematic for all $k\geq 4$ and no nematic-to-isotropic transition exists at high densities, similarly to the BL solution drs11 .
@@
The values of the critical activities, densities, and entropies for $4\leq k\leq 15$, are summarized in Table 1. We notice that the critical densities and entropies are monotonically decreasing functions of $k$, but the critical activities show a minimum at $k=10$, followed by a slow increase. We may compare the critical parameters with the ones found for the model on the four coordinated Bethe lattice in drs11 , which are:
@@
We found that, as expected, the absolute values of the relative differences between the values of these parameters on the two lattices are decreasing functions of $k$, starting with values of the order of 20% for $k=4$ and ending with values around 0.5% or lower.
@@
The entropy per lattice site, also as a function of the density of sites occupied by monomers, is depicted in Fig. 5. It obviously starts at zero for $\rho=0$, reaches a maximum at an intermediate density, which grows monotonically with $k$, and ends at the values already obtained in Ref. rso22 in the full lattice limit.
@@
As expected, when we compare the results found on the Husimi lattice with the behavior of the model on the Bethe lattice drs11 , in general, the results on the Husimi lattice are somewhat closer to the estimates provided by numerical simulations of the system on the square lattice. For instance, as discussed in the Introduction, on the square lattice the nematic order parameter vanishes in the full lattice limit, while both in the Bethe and Husimi lattices it does not vanish. However, as demonstrated in rso22 , $\Psi$ is always smaller on the Husimi lattice than on the Bethe lattice; see table III of Ref. rso22 . It is also noteworthy in such a table, as it was seen in the comparisons between critical parameters on the Bethe and Husimi lattices above, that the difference between the order parameters on the two lattices becomes smaller as $k$ increases, since the ratio between the size of the rods and the size of the elementary squares of the lattice grows. This effect is observed in all comparisons between both lattices, being clearly visible, e.g., in the full lattice entropies presented in table III of Ref. rso22 .
The values of the critical density on the HL as a function of the size of the rod $k$ are compared with those for the BL and for the square lattice in Fig. 6. The latter ones were obtained by Matoz-Fernandez et al. mf08 ; mf08p , according to which $\rho_{c}(k)\sim k^{-1}$ and, moreover, $\rho_{c}(10)\approx 0.502$ for rigid rods on the square lattice, such that $\rho_{c}(k)\simeq 5.02/k$. We notice that, for $k=7$ this expression gives $\rho_{c}(7)\approx 0.717$, which is not so different from the estimate reported by Kundu et al. k13 : $\rho_{c}(7)\approx 0.745$. Recalling that results on the treelike lattices are mean-field approximations of the model on the square lattice, we expect them to underestimate the critical density. This is indeed confirmed in Fig. 6, where we may see that the critical densities on the square HL are larger than those found on the Bethe lattice. Of course, the mean-field results are still much lower than the available estimates for the square lattice ones.
Figure 6: Critical densities on Bethe and square HL, as functions of $k$. Circles correspond to the four coordinated Bethe lattice results drs11 , squares to Husimi lattice results and the blue line are simulational estimates for the model on the square lattice mf08 ; mf08p
IV Results for rods on the triangular HL
IV.1 Isotropic-nematic transition
Similarly to the results above for the square HL (and also to those for the BL drs11 ), the $k$-mers never present nematic order in the triangular HL for $k=2$ and $3$, whereas an isotropic-nematic transition is found for $k\geq 4$. Such a transition is always discontinuous in the triangular HL case, in consonance with the BL approach for coordination $q=6$ drs11 . This is shown in Fig. 7, where one sees that both the monomer density $\rho$ and the order parameter $\Psi$ present a discontinuous behavior as $z$ increases.
@@ IV.1 Isotropic-nematic transition
Noteworthy, the scenario above is very similar to that found for rods on the BL with $q=6$ drs11 . It is important to remark, however, that the coexistence points were estimated in Ref. drs11 by writing down the entropy [Eq. (32) there] as a function of $\rho$ and $\Psi$, and then defining the “critical density” $\rho_{c}$ as the value for which $s(\Psi,\rho=\rho_{c})$ has two peaks of equal height  [^1] When comparing the results here with those in Ref. drs11 , it is important to bear in mind that the activity there (let us call it $\tilde{z}$) is for rod, while here $z$ is for a single monomer, such that $z=\tilde{z}^{1/k}$.. We have revisited such calculations for the six-coordinate BL here, but determining the coexistence points through the free energy — i.e., by requiring that $\phi_{b}^{(I)}=\phi_{b}^{(N)}$ —, which gives results slightly different from those reported in drs11 .
@@
Figures 8(a)-(d) present the outcomes from such BL solution, comparing them with those for the triangular HL. In general, all quantities have larger values on the BL, but the same qualitative behavior. For instance, $z^{*}$ has also a non-monotonic variation with $k$ (with a minimum at $k=13$) in the BL case. The same non-monotonic behavior is observed in the spinodals of both the triangular HL (see Tab. 2) and the BL. In fact, for the BL the spinodal of the isotropic phase can be exactly determined, being drs11
@@
Moreover, although the transitions are discontinuous in both the BL and triangular HL solutions, the nematic order parameters at coexistence are much smaller in the HL case, especially for the smaller $k$’s. These results demonstrate that, by improving the approximation, one obtains a behavior closer to a continuous transition, as expected for the triangular lattice mf08 ; mf08p . Curiously, however, the coexistence densities do not follow this trend. In fact, from Monte Carlo simulations of rigid rods on the triangular lattice and some theoretical approaches, Matoz-Fernandez et al. mf08 ; mf08p found that the continuous isotropic-nematic transitions occur at critical densities given by $\rho_{c}(k)\simeq 5.31/k$, for $k\geq 7$. As it may be observed in Figs. 8(c) and 8(d), these critical densities are much larger than the coexistence ones found here. Moreover, both $\rho^{*(I)}(k)$ and $\rho^{*(N)}(k)$, are more distant from $\rho_{c}(k)$ in the HL case.
IV.2 The full-packing limit
@@
The entropies of the isotropic and nematic phases are presented in Table 3, for $2\leq k\leq 7$. In agreement with the results from the previous subsection, only the isotropic phase is stable in the full lattice case for dimers and trimmers. For larger $k$’s, however, the entropy of the isotropic phase is negative, while that for the nematic phase is positive. Therefore, for $k\geq 4$, the nematic phase is the single one stable for large $z$ and this persists until the fully occupancy limit. Hence, no nematic-isotropic transition close to full-packing — as observed for rigid rods on the square lattice k12 — occurs in the triangular HL. The nematic order parameters are also shown in Tab. 3, which are close, but smaller than 1.
For comparison, Table 3 presents also the entropies and order parameters for rigid rods placed on the Bethe lattice (BL) with coordination $q=6$, in the full lattice limit, as obtained in Ref. drs11 . As in the HL case, the stable phase is isotropic for $k\leq 3$ and nematic for $k\geq 4$. Interestingly, the entropies of these phases are very similar in the BL and HL approximations. For instance, they differ by less than 2% in the case of dimers, for which they are very close also to the exact value on the triangular lattice: $s(3)=0.428594537\ldots$ (see, e.g., Ref. Wu06 and references therein). Remarkably, the difference between the HL and the exact result for dimers is only 0.7%. This is certainly a consequence of the fact that the dimer-dimer correlations on the triangular lattice are short ranged Fendley02 , so that the weakening of correlations caused by the treelike structure of the HL has a small effect on the properties of this model. Also, this suggests that generalized triangular HLs — built, e.g., with hexagonal triangular lattice cells, in the same spirit of the generalized square HLs considered in Refs. ro21 ; rso22 — may give the exact solution of the dimer problem for finite (and perhaps small) cells.
@@
As demonstrated in Table III of Ref. rso22 , which shows the same results as presented in Table 3 here, but for the BL with $q=4$ and for the square HL — the rods’ entropies in these four-coordinated lattices are smaller than those in the six-coordinated cases analyzed here, as expected. Moreover, while the entropy for a given $k$ converges from below in $q=4$ case (see Tabs. III and V, and Fig. 2 in Ref. rso22 ), for the lattices considered here (with $q=6$) it converges from above to the asymptotic (triangular lattice) behavior. Namely, the number of possible $k$-mers’ configurations decreases as one goes from the BL ($q=6$) to the triangular HL and, then, to the triangular lattice; while the opposite behavior is observed for the lattices with $q=4$.
V Discussion
@@
We have studied the model of $k$-mers, rigid rods composed by $k$ monomers placed on lattice sites with $k-1$ edges of the lattice between them, all aligned in the same direction, with excluded volume interactions only; that is, only configurations of the rods where each lattice site is occupied by at most one monomer are accepted. We obtained the solution of this model in the grand-canonical ensemble in the central region of two treelike lattices built with polygons (Husimi trees): the tree built with squares, so that two squares meet at each lattice site (ramification of squares equal to one and coordination number $q=4$); and the one built with triangles, with three of them meeting at each site (ramification of triangles equal to two and $q=6$). The exact solution in the core of these trees corresponds to a mean-field approximation of the model on the square and triangular lattice, respectively. Due to the existence of loops in the Husime trees, these calculations may be considered to be higher order mean-field approximations to the problem as compared to the Bethe lattice solution, which was obtained in Ref. drs11 .
@@
On the Husimi lattice built with squares, the rods display a continuous isotropic-nematic transition, as it is also the case on the four-coordinated Bethe lattice and on numerical simulations on the square lattice. In contrast, a discontinuous transition was found for the model on the Husimi lattice built with triangles, as happens also on the six-coordinated Bethe lattice, while numerical simulations on the triangular lattice suggest a continuous transition in the class of the 3-state Potts model. Therefore, in this case, the Husimi lattice solution also does not lead to the order of the transition obtained using numerical simulations. It is worthy remarking, however, that mean-field approaches to the 3-state Potts model are known to display a discontinuous transition w82 , which suggests that the same thing can be happen in the rod model. Namely, the discontinuous transition found for $k$-mers in the mean-field solution on the triangular Husimi lattice may be viewed as an indication that this transition is indeed in the 3-state Potts class in the triangular lattice. We notice, for instance, that a similar scenario was recently found for an associating lattice gas model, which exhibits a continuous transition in the 3-state Potts class between a fluid a high-density liquid phase, as revealed by Monte Carlo simulations and transfer matrix calculations furlanALG ; ibagon21 , but it was found to be discontinuous in Husimi lattice solutions tiago10 ; furlanALG . Furthermore, we recall that numerical simulations of the rod model on the cubic lattice vdr17 suggested the isotropic-nematic transition to be continuous for $k\geq 7$, but the possibility of a weak discontinuous first-order transition, as is found for the 3-state Potts model on this lattice w82 , was not ruled out, although it is not clearly supported by the simulational results. Once again, the isotropic-nematic transitions found in the early studies via mean-field approximations of the three-dimensional models mentioned before o49 ; f56 ; z63 were also discontinuous.
@@
If we consider the critical densities at which the transition happen in our calculations for the square case, they are always between the ones found on the Bethe lattice and the estimates coming from the simulations. The same is true for the entropies in the full lattice limit, so that for these parameters the Husimi lattice results are closer to the ones on the corresponding Bravais lattices than the ones coming from the solution on the Bethe lattice, as expected. On the other hand, the coexistence densities obtained in the triangular case display an opposite trend, suggesting that they converge non-monotonically to the triangular lattice results as the mean-field approximation is improved. Another point where there was no improvement in the Husimi lattice solutions when compared to the ones on the Bethe lattice was the minimum value of $k_{min}$ for which the phase transition was observed, which is equal to four in all cases, as opposed to $k_{min}=7$ for the square and triangular lattices. Calculations in the full lattice limit rso22 on treelike lattices built with larger square cells, such as the Kobayashi-Susuki lattice ks93 , indicate that larger values of $k_{min}$ are found when the size of the cells grows. Therefore, it is interesting to extend this study to include vacancies in the configurations of the model on such generalized Husimi lattices to find out the complete thermodynamic behavior of rigid rods on them. In a similar vein, investigating the $k$-mers on such lattices built with triangular lattice cells can be also very important to verify whether the isotropic-nematic transition turns out to be continuous at some point, as well as to estimate the entropies in the full-packing limit, since it seems that they have never been calculated in the literature for the triangular lattice, beyond the dimer case. We are currently working in these projects.

### v2

@@ 3 Results for rods on the square HL
This is indeed confirmed in Fig. 4, where the order parameter is shown as a function of the density of sites occupied by the rods on the lattice, for different values of $k$. As $k$ grows, $\Psi$ approaches unity in the limit of full occupancy (corresponding to $z\rightarrow\infty$ and, thus, $\rho=1$), but it does not reach it, as already demonstrated in Ref. [27]. Therefore, in such a limit the stable phase is nematic for all $k\geq 4$ and no nematic-to-isotropic transition exists at high densities, similarly to the BL solution [37].
@@
The values of the critical activities, densities, and entropies for $4\leq k\leq 15$, are summarized in Table 1. We notice that the critical densities and entropies are monotonically decreasing functions of $k$, but the critical activities show a minimum at $k=10$, followed by a slow increase. We may compare the critical parameters with the ones found for the model on the four coordinated Bethe lattice in [37], which are:
@@
We recall that the Bethe and Husimi lattice solutions of a model are mean-field approximations, and the second one is of higher order than the first. It is thus interesting to compare the results of both calculations and eventual other available estimates, as the ones provided by simulations. We found that, as expected, the absolute values of the relative differences between the values of these parameters on the two lattices are decreasing functions of $k$, starting with values of the order of 20% for $k=4$ and ending with values around 0.5% or lower.
@@
The entropy per lattice site, also as a function of the density of sites occupied by monomers, is depicted in Fig. 5. It obviously starts at zero for $\rho=0$, reaches a maximum at an intermediate density, which grows monotonically with $k$, and ends at the values already obtained in Ref. [27] in the full lattice limit.
@@
As expected, when we compare the results found on the Husimi lattice with the behavior of the model on the Bethe lattice [37], in general, the results on the Husimi lattice are somewhat closer to the estimates provided by numerical simulations of the system on the square lattice. For instance, as discussed in the Introduction, on the square lattice the nematic order parameter vanishes in the full lattice limit, while both in the Bethe and Husimi lattices it does not vanish. However, as demonstrated in [27], $\Psi$ is always smaller on the Husimi lattice than on the Bethe lattice; see table III of Ref. [27]. It is also noteworthy in such a table, as it was seen in the comparisons between critical parameters on the Bethe and Husimi lattices above, that the difference between the order parameters on the two lattices becomes smaller as $k$ increases, since the ratio between the size of the rods and the size of the elementary squares of the lattice grows. This effect is observed in all comparisons between both lattices, being clearly visible, e.g., in the full lattice entropies presented in table III of Ref. [27].
The values of the critical density on the HL as a function of the size of the rod $k$ are compared with those for the BL and for the square lattice in Fig. 6. The latter ones were obtained by Matoz-Fernandez et al. [29, 30], according to which $\rho_{c}(k)\sim k^{-1}$ and, moreover, $\rho_{c}(10)\approx 0.502$ for rigid rods on the square lattice, such that $\rho_{c}(k)\simeq 5.02/k$. We notice that, for $k=7$ this expression gives $\rho_{c}(7)\approx 0.717$, which is not so different from the estimate reported by Kundu et al. [35]: $\rho_{c}(7)\approx 0.745$. Recalling that results on these hierarchical lattices are mean-field approximations of the model on the square lattice, we expect them to underestimate the critical density. This is indeed confirmed in Fig. 6, where we may see that the critical densities on the square HL are larger than those found on the Bethe lattice. Of course, the mean-field results are still much lower than the available estimates for the square lattice ones.
Figure 6: Critical densities on Bethe and square HL, as functions of $k$. Circles correspond to the four coordinated Bethe lattice results [37], squares to Husimi lattice results and the blue line are simulational estimates for the model on the square lattice [29, 30]
4 Results for rods on the triangular HL
4.1 Isotropic-nematic transition
Similarly to the results above for the square HL (and also to those for the BL [37]), the $k$-mers never present nematic order in the triangular HL for $k=2$ and $3$, whereas an isotropic-nematic transition is found for $k\geq 4$. Such a transition is always discontinuous in the triangular HL case, in consonance with the BL approach for coordination $q=6$ [37]. This is shown in Fig. 7, where one sees that both the monomer density $\rho$ and the order parameter $\Psi$ present a discontinuous behavior as $z$ increases.
@@ 4.1 Isotropic-nematic transition
Noteworthy, the scenario above is very similar to that found for rods on the BL with $q=6$ [37]. It is important to remark, however, that the coexistence points were estimated in Ref. [37] by writing down the entropy [Eq. (32) there] as a function of $\rho$ and $\Psi$, and then defining the “critical density” $\rho_{c}$ as the value for which $s(\Psi,\rho=\rho_{c})$ has two peaks of equal height  [^2] When comparing the results here with those in Ref. [37], it is important to bear in mind that the activity there (let us call it $\tilde{z}$) is for rod, while here $z$ is for a single monomer, such that $z=\tilde{z}^{1/k}$.. We have revisited such calculations for the six-coordinate BL here, but determining the coexistence points through the free energy — i.e., by requiring that $\phi_{b}^{(I)}=\phi_{b}^{(N)}$ —, which gives results slightly different from those reported in [37].
@@
Figures 8(a)-(d) present the outcomes from such BL solution, comparing them with those for the triangular HL. In general, all quantities have larger values on the BL, but the same qualitative behavior. For instance, $z^{*}$ has also a non-monotonic variation with $k$ (with a minimum at $k=13$) in the BL case. The same non-monotonic behavior is observed in the spinodals of both the triangular HL (see Tab. 2) and the BL. In fact, for the BL the spinodal of the isotropic phase can be exactly determined, being [37]
@@
Moreover, although the transitions are discontinuous in both the BL and triangular HL solutions, the nematic order parameters at coexistence are much smaller in the HL case, especially for the smaller $k$’s. These results demonstrate that, by improving the approximation, one obtains a behavior closer to a continuous transition, as expected for the triangular lattice [29, 30]. Curiously, however, the coexistence densities do not follow this trend. In fact, from Monte Carlo simulations of rigid rods on the triangular lattice and some theoretical approaches, Matoz-Fernandez et al. [29, 30] found that the continuous isotropic-nematic transitions occur at critical densities given by $\rho_{c}(k)\simeq 5.31/k$, for $k\geq 7$. As it may be observed in Figs. 8(c) and 8(d), these critical densities are much larger than the coexistence ones found here. Moreover, both $\rho^{*(I)}(k)$ and $\rho^{*(N)}(k)$, are more distant from $\rho_{c}(k)$ in the HL case.
4.2 The full-packing limit
@@
The entropies of the isotropic and nematic phases are presented in Table 3, for $2\leq k\leq 7$. In agreement with the results from the previous subsection, only the isotropic phase is stable in the full lattice case for dimers and trimers. For larger $k$’s, however, the entropy of the isotropic phase is negative, indicating that this phase is unstable, while that for the nematic phase it is positive. Therefore, for $k\geq 4$, the nematic phase is the single one stable for large $z$ and this persists until the fully occupancy limit. Hence, no nematic-isotropic transition close to full-packing — as observed for rigid rods on the square lattice [34] — occurs in the triangular HL. The nematic order parameters are also shown in Tab. 3, which are close, but smaller than 1. In addition, in agreement with previous results for the Bethe lattice (BL) [37] and the square HL [27], they get close to $\Psi=1$ as $k$ increases.
For comparison, Table 3 presents also the entropies and order parameters for rigid rods placed on the BL with coordination $q=6$, in the full lattice limit, as obtained in Ref. [37]. As in the HL case, the stable phase is isotropic for $k\leq 3$ and nematic for $k\geq 4$. Interestingly, the entropies of these phases are very similar in the BL and HL approximations. For instance, they differ by less than 2% in the case of dimers, for which they are very close also to the exact value on the triangular lattice: $s_{2}^{(tri)}=0.428594537\ldots$ (see, e.g., Ref. [22] and references therein). Remarkably, the difference between the HL and the exact result for dimers is only 0.7%. This is certainly a consequence of the fact that the dimer-dimer correlations on the triangular lattice are short ranged [21], so that the weakening of correlations caused by the hierarchical structure of the HL has a small effect on the properties of this model. Also, this suggests that generalized triangular HLs — built, e.g., with hexagonal triangular lattice cells, in the same spirit of the generalized square HLs considered in Refs. [46, 27] — may give the exact solution of the dimer problem for finite (and perhaps small) cells.
@@
As demonstrated in Table III of Ref. [27], which shows the same results as presented in Table 3 here, but for the BL with $q=4$ and for the square HL — the rods’ entropies in these four-coordinated lattices are smaller than those in the six-coordinated cases analyzed here, as expected. Moreover, while the entropy for a given $k$ converges from below in $q=4$ case (see Tabs. III and V, and Fig. 2 in Ref. [27]), for the lattices considered here (with $q=6$) it converges from above to the asymptotic (triangular lattice) behavior. Namely, the number of possible $k$-mers’ configurations decreases as one goes from the BL ($q=6$) to the triangular HL and, then, to the triangular lattice; while the opposite behavior is observed for the lattices with $q=4$.
5 Discussion
@@
We have studied the model of $k$-mers, rigid rods composed by $k$ monomers placed on lattice sites with $k-1$ edges of the lattice between them, all aligned in the same direction, with excluded volume interactions only; that is, only configurations of the rods where each lattice site is occupied by at most one monomer are accepted. We obtained the solution of this model in the grand-canonical ensemble in the central region of two hierarchical lattices built with polygons (Husimi cactuses): the cactus built with squares, so that two squares meet at each lattice site (ramification of squares equal to one and coordination number $q=4$); and the one built with triangles, with three of them meeting at each site (ramification of triangles equal to two and $q=6$). The exact solution in the core of these cactuses corresponds to a mean-field approximation of the model on the square and triangular lattice, respectively. Due to the existence of loops in the Husimi lattices, these calculations may be considered to be higher order mean-field approximations to the problem as compared to the Bethe lattice solution, which was obtained in Ref. [37].
@@
On the Husimi lattice built with squares, the rods display a continuous isotropic-nematic transition, as it is also the case on the four-coordinated Bethe lattice and on numerical simulations on the square lattice. In contrast, a discontinuous transition was found for the model on the Husimi lattice built with triangles, as happens also on the six-coordinated Bethe lattice, while numerical simulations on the triangular lattice suggest a continuous transition in the class of the 3-state Potts model. Therefore, in this case, the Husimi lattice solution also does not lead to the order of the transition obtained using numerical simulations. It is worthy remarking, however, that mean-field approaches to the 3-state Potts model are known to display a discontinuous transition [48], which suggests that the same thing can be happen in the rod model. So, in this sense, the discontinuous transition found for k-mers in the mean-field solution on the triangular Husimi lattice is consistent with the 3-state Potts scenario for the triangular lattice. We notice, for instance, that a similar scenario was recently found for an associating lattice gas model, which exhibits a continuous transition in the 3-state Potts class between a fluid and a high-density liquid phase, as revealed by Monte Carlo simulations and transfer matrix calculations [45, 50], but it was found to be discontinuous in Husimi lattice solutions [51, 45]. Furthermore, we recall that numerical simulations of the rod model on the cubic lattice [47] suggested the isotropic-nematic transition to be continuous for $k\geq 7$, but the possibility of a weak discontinuous first-order transition, as is found for the 3-state Potts model on this lattice [48], was not ruled out, although it is not clearly supported by the simulational results. Once again, the isotropic-nematic transitions found in the early studies via mean-field approximations of the three-dimensional models mentioned before [11, 12, 13] were also discontinuous.
@@
If we consider the critical densities at which the transition happen in our calculations for the square case, they are always between the ones found on the Bethe lattice and the estimates coming from the simulations. The same is true for the entropies in the full lattice limit, so that for these parameters the Husimi lattice results are closer to the ones on the corresponding Bravais lattices than the ones coming from the solution on the Bethe lattice, as expected. On the other hand, the coexistence densities obtained in the triangular case display an opposite trend, suggesting that they converge non-monotonically to the triangular lattice results as the mean-field approximation is improved. Another point where there was no improvement in the Husimi lattice solutions when compared to the ones on the Bethe lattice was the minimum value of $k_{min}$ for which the phase transition was observed, which is equal to four in all cases, as opposed to $k_{min}=7$ for the square and triangular lattices. Calculations in the full lattice limit [27] on hierarchical lattices built with larger square cells, such as the Kobayashi-Susuki lattice [49], indicate that larger values of $k_{min}$ are found when the size of the cells grows. Therefore, it is interesting to extend this study to include vacancies in the configurations of the model on such generalized Husimi lattices to find out the complete thermodynamic behavior of rigid rods on them. In a similar vein, investigating the $k$-mers on such lattices built with triangular lattice cells can be also very important to verify whether the isotropic-nematic transition turns out to be continuous at some point, as well as to estimate the entropies in the full-packing limit, since it seems that they have never been calculated in the literature for the triangular lattice, beyond the dimer case. We are currently working in these projects.